### v1

# Hydrostatic mass of galaxy clusters within some theories of gravity

Abstract
The mass of galaxy clusters (GCs) can be determined by calculating the hydrostatic equilibrium equation. In this work, we revisit the derivation of hydrostatic mass of GCs within modified emergent Newtonian gravity (MENG) with generalized uncertainty principle (GUP) correction, Eddington-inspired Born-Infeld (EiBI) theory, and beyond Horndeski gravity (BHG). We apply the formulations on the masses of 10 GCs. We compare our results with the Newtonian mass of GCs and their baryonic masses $M_{bar}$. We find that all formulations could match with the Newtonian mass and the impact of the modified theories of gravity used in this work can be neglected. The noteworthy impact for MENG starts if we set $\beta_{0}=-1.656\times 10^{110}$, $\kappa=5\times 10^{40}$ for EiBI theory, and $\Upsilon=-0,1655\times 10^{69}$ for BHG. In contrast, for the comparison with baryonic mass, only the MENG formulation produces the best linear fit with the slope $\frac{M_{MENG}}{M_{bar}}=0.755\pm 0.061$. It means that in this work, MENG is the best theory of gravity that alleviates the mass discrepancy of hydrostatic mass and baryonic mass of GCs.
[…]
1 Introduction
Galaxy clusters (GCs) are the largest virialized astronomical structures in the universe [1]. They are generally defined as self-gravitating systems of galaxies that may be identified in observations as the systems exhibiting one or more of the following features: a significantly higher concentration of galaxies relative to the average distribution of the large-scale structure of the universe; massive concentration(s) of X-ray emitting gas, often associated with the Sunyaev-Zel’dovich (SZ) effect on the Cosmic Microwave Background Radiation; and a non-random gravitational lensing shear map of background galaxies [2]. As a relatively large structure, GCs are magnificent laboratories for constraining fundamental physics, such as bounding the neutrino mass, graviton mass, or limits on primordial non-Gaussianity. GCs also have an important role in cosmology due to the many discoveries based on optical, X-ray, and microwave surveys. [3].
GCs are full of hot, diffuse plasma, with typical temperatures $T\sim 7\times 10^{7}$ K and electron densities $n_{e}\sim 10^{-3}$ cm-3, This intraclustermedium leads to emission of copious X-ray, with typical cluster X-ray luminosities $L_{x}\sim 10^{43}-10^{45}$ ergs s-1. The total mass of intracluster gas generally transcends the mass of all of the galaxies in a rich cluster. The gas might be almost hydrostatic. From the distribution of the intracluster plasma, one can obtain the total gravitational mass of the cluster and its distribution [4].
Traditionally, mass of GCs can be determined using three methods, i.e. velocity dispersion, X-ray profiles or SZ observations assuming hydrostatic equilibrium, and gravitational lensing. Apart from gravitational lensing, all those methods produce formulation of GC’s mass within Newtonian approximation [3]. Review on all mentioned methods can be found in Refs. [5, 6, 7, 8, 9, 10, 11, 12, 13, 14].
Bambi [15] profoundly worked on the problem of GC’s mass. Starting from an assumption that dark energy (DE) behaves like cosmological constant $\Lambda$ on a relatively small scale, he discussed the effects of a non-zero cosmological constant in the Newtonian limit. As a result, he derived the general relativity (GR) correction for the mass of GCs, which leads to such an effective mass of GCs, i.e. $M_{0}-\frac{8}{3}\pi r^{3}\rho_{\Lambda}$, where $M_{0}$ is the Newtonian mass of GCs, and $\rho_{\Lambda}$ is the energy density associated with the cosmological constant $\Lambda=8\pi G\rho_{\Lambda}$, and $G$ denotes the universal gravitational constant. It is worthy to note that the masses of GCs obtained from velocity dispersion and X-ray profile methods are the effective mass, while the masses obtained from gravitational lensing are the pure Newtonian masses of GCs. Furthermore, even though gravitational lensing is a GR phenomenon, Bambi argued that the gravitational lensing measurement of GC’s mass is not affected by $\Lambda$. Based on Bambi’s formulation, Desai [16] calculated masses of several GCs. He found that the impact of general relativistic corrections to hydrostatic masses of galaxy clusters using the Kottler metric (i.e. the metric used in Bambi’s work [15]) is negligible.
Gupta & Desai [3] proposed other formulation to find GR-type correction on GC’s mass, i.e. using Tolman-Oppenheimer-Volkoff (TOV) equation. They then applied the GC’s mass formulation they have obtained to calculate the hydrostatic masses of twelve GCs. It is important to keep in mind that TOV equation is the GR form of Newtonian hydrostatic equilibrium equation. The differences between the TOV based mass and the Newtonian mass are corresponding to $\mathcal{O}(10^{-5})$, so they can be neglected. It is concluded that for all practical purposes, we can safely use the Newtonian equation of hydrostatic equilibrium for any dynamical analysis with GCs.
Another attempt to obtain mass of GCs was achieved by Apryandi et. al. [17]. Using generalized uncertainty principle (GUP) and following Verlinde’s emergent gravity or entropic gravity procedure [18] to get Newtonian gravity formula, they derived the modified emergent Newtonian gravity (MENG). The formalism was then used to derive a formula to calculate mass of GCs. Unfortunately, in their work, they have not yet applied their formula to calculate GCs’ masses, like what has been done by Desai [16] and Gupta & Desai [3].
Basically, Verlinde’s theory of emergent gravity describes gravity as an emergent phenomenon rather than a fundamental force. One can impose this reasoning on de Sitter space to find that gravity behaving differently on galaxy and GC scales. This excess gravity might offer an alternative to dark matter (DM) [19]. The emergent gravity paradigm is such an attempt to obtain and interpret the field equations of gravitational theories in a thermodynamic context [20]. Please see Refs. [18, 19, 20] for further discussions on emergent gravity.
Unfortunately, Tamosiunas et al. [19] found that GR is preferred over emergent gravity in all tested datasets of galaxy and GCs. They discovered that the simultaneous emergent gravity fits of the X-ray and weak lensing datasets are significantly worse than those provided by GR with cold dark matter. For the Coma cluster, the predictions from emergent gravity and GR agree in the range of 250--700 kpc. In contrast, at around 1 Mpc scales, emergent gravity total mass predictions are larger by a factor of 2. For the cluster stack the predictions are only in good agreement at around the 1--2 Mpc scales, while for the radius $r\gtrsim$ 10 Mpc, emergent gravity is in strong tension with the data. It is important to keep in mind that GUP’s effect is not included in the work of Tamosiunas et al. Although controversial [21, 22, 23], Verlinde’s theory of emergent gravity has good agreement with some observations (please see Refs. [24, 25, 26, 27]).
Other than the work of Apryandi et al. [17], some attempts to involve GUP in Verlinde’s theory have been achieved. Feng et al. [28] found that the GUP corrected entropic force is related to the properties of the black holes and the Planck length. Furthermore, based on the GUP corrected entropic force, they derive the modified Einstein equations and the modified Friedmann equation. Kibaroğlu [29] obtained generalized Einstein equations with cosmological constant by including GUP in emergent gravity and using the modified Unruh temperature. On the other side, motivated by the existence of some white dwarfs which are more massive than the Chandrasekhar mass limit, Belfaqih et al. [30] included GUP in Verlinde’s theory of emergent gravity to find mass and radius relation of white dwarfs. Jizba et al. [31] unified GUP and Tsallis thermostatistics, and then applied their achievement to Verlinde’s theory. A discussion on Tsallis thermostatistics can be found in Ref. [32].
On the other hand, beyond GR theories or modified theories of gravity are also important to be considered in the studies of objects in the cosmos. For decades, people have been witnessing the evidence that, if gravity is described by GR, there should be a considerable amount of DM in galaxies and GCs. Furthermore, DE has become a problem that needs to be explained as the expansion of the Universe is still accelerated. However, if GR is correct, it must be around 96 % of the Universe should be in the form of energy densities that do not have any electromagnetic interaction [33]. In the context of GCs, observations of GCs suggest that they contain fewer baryons (gas plus stars) than the cosmic baryon fraction. This so-called “missing baryon” is surprising for the most massive GCs, which are expected to be representative of the cosmic matter content of the Universe (baryons and DM) [34]. These facts have led us to a speculation on the possibility that GR may not, apparently, be the correct theory of gravity to describe the Universe on the largest scales. Our dark universe may be just another sign that we have to go beyond Einstein’s GR [33].
An example of modified theory of gravity is Eddington-inspired Born-Infeld (EiBI) theory. Santos & Santos [35] introduced geometrical mass of GCs within EiBI theory. The main problem they pointed out is the mass discrepancy in GCs, which is usually explained by a postulate of the existence of a cold pressureless DM. Unfortunately, the nature of DM is still unknown and the only existing evidence for its presence is gravitational. Thus, another possible way to overcome the problem of discrepancy of the GC’s mass is modifying the theory of gravity. In their work, they derived the virial theorem in EiBI theory, and showed that the geometrical mass, which is a mathematical artifact in the theorem, may account for the virial mass discrepancy in GCs.
Another example of modified theory of gravity is beyond Horndeski gravity (BHG), which is a modified gravity containing the screening mechanism if $R<<r_{v}$, where $R$ is the radius of the corresponding astrophysical body, and $r_{v}$ is the Vainshtein radius. Here $r_{v}$ is defined as the transition between the screened and unscreened regimes inside astrophysical bodies [36]. The Vainshtein mechanism is discussed in Refs. [37, 38] and the references therein. Sakstein et. al. [39] used GCs to obtain new constraints on BHG. In their work, they presented the so-called thermal mass, which was previously introduced by Terukina et.al. [40]. In Ref. [39], the thermal mass is equivalent to weak-lensing mass of GCs, and found by solving the hydrostatic equilibrium equation for the thermal part i.e. the part deduced from X-ray temperature profiles.
Motivated by the previous studies, in this work we calculate hydrostatic mass of several GCs presented in Ref. [3] using MENG formulation derived by Apryandi et. al. [17]. Beside that, we also derive the formulation for hydrostatic mass of GCs within EiBI theory and BHG, and then obtain EiBI and BHG version of numerical calculation on hydrostatic mass of the GCs which is previously numerically calculated using MENG. Finally, the results are compared with the corresponding Newtonian values and with the baryonic mass $M_{bar}$ of the GCs presented in Ref. [41]. We expect that the baryonic mass and the hydrostatic mass would essentially agree if the free parameters of MENG, EiBI, and BHG were correctly represented by our formulation.
This paper is organized as follows. In section 2, we explain the construction of formalism we use, where the MENG formulation derived in Ref. [17] is revisited, and the derivations of hydrostatic mass of GCs within EiBI theory and BHG are presented. In section 3, we present and discuss the numerical results of GCs masses calculation. Lastly, section 4 contains concluding remarks of this work.
2 Formalisms
To make the discussion self-contained, in subsection 2.1, we revisit the hydrostatic mass of GCs within MENG formalism derived by Apryandi et al. [17]; in subsection 2.2 we derive hydrostatic mass of GCs within EiBI theory; while derivation for hydrostatic mass of GCs within BHG is given in subsection 2.3.
2.1 Hydrostatic mass of GCs within MENG
In this subsection, we start with the GUP. It is worthy to keep in mind the ordinary unmodified Heisenberg uncertainty, i.e. $\Delta x\Delta p\geq\frac{\hbar}{2}$. Here we use the quadratic GUP, which reads
$\Delta x\Delta p\geq\frac{\hbar}{2}\left(1+\beta p^{2}\right),$ (1)
where $\beta=\frac{\beta_{0}l_{p}^{2}}{\hbar^{2}}$, with $\beta_{0}$ is a free dimensionless parameter of GUP, and Planck length $l_{p}=\sqrt{\frac{G\hbar}{c^{3}}}\approx$ 10-35 m. Moreover, the modified minimum volume of phase-space writes
$\Delta V_{GUP}=\hbar^{3}\left(1+\beta p^{2}\right)^{3}.$ (2)
Due to Eq. (2), the volume invariance of phase-space has to be transformed, i.e.
$\int\frac{d^{3}x\>d^{3}p}{\hbar^{3}}\longrightarrow\int\frac{d^{3}x\>d^{3}p}{\hbar^{3}\left(1+\beta p^{2}\right)^{3}}.$ (3)
Based on the transformed volume invariance showed by Eq. (3), we can write the expression of energy surface equation, i.e.
$\sum E=\frac{1}{h^{3N}}\int\frac{d^{3}x_{1}...d^{3}x_{N}\>d^{3}p_{1}...d^{3}p_{N}}{(1+\beta p^{2})^{3N}}.$ (4)
Here $N$ denotes the amount of particles which is related to the equation of state (EOS) of ideal gas, i.e. $PV=Nk_{b}T$, where $P$ is pressure, $V$ is volume, $k_{b}$ is Boltzmann constant, and $T$ is absolute temperature. The EOS of ideal gas becomes very important in this study since we follow the procedure in Ref. [3] in using the EOS of ideal gas as the EOS of GCs.
By applying $p^{2}=2m_{b}E$, where $m_{b}$ denotes the mass of an individual atom of the ideal gas; and the entropy $S=k_{b}\ln(\sum E)$, it is easy to show that
$S=Nk_{b}\left[\frac{5}{2}+\ln\left\{\left(\frac{V}{N\hbar^{3}(1+2\beta m_{b}E)^{3}}\right)\left(\frac{4\pi m_{b}E}{3N}\right)^{3/2}\right\}\right].$ (5)
The absolute temperature $T=\frac{\partial E}{\partial S}$ now becomes
$T=\left(\frac{1+2\beta m_{b}E}{1-2\beta m_{b}E}\right)\left(\frac{2E}{3k_{b}N}\right);$ (6)
while the pressure $P=-\frac{\partial E}{\partial V}$ writes
$P=\left(\frac{1+2\beta m_{b}E}{1-2\beta m_{b}E}\right)\frac{2E}{3V}.$ (7)
On the other hand, a system of bits on the holographic surface satisfies the ideal gas distribution. The amount of information or the amount of bits on the holographic surface is given by [18]
$N=\frac{Ac^{3}}{G\hbar},$ (8)
where $A$ denotes the surface area, and $c$ denotes the speed of light. Considering Bekenstein-Hawking entropy [42, 43] $S_{BH}=\frac{k_{b}A}{4l_{p}^{2}}$, Eq. (8) now reads
$N=\frac{4S_{BH}}{k_{b}}.$ (9)
After involving GUP, Eq. (9) is modified to be
$\mathcal{N}=\frac{A}{l_{p}^{2}}-\frac{\pi\beta_{0}}{4}\ln\left(\frac{A}{4l_{p}^{2}}\right);$ (10)
while Bekenstein-Hawking entropy is modified to be
$\mathcal{S}_{BH}=k_{b}\frac{A}{4l_{p}^{2}}-\frac{k_{b}\pi\beta_{0}}{16}\ln\left(\frac{A}{4l_{p}^{2}}\right).$ (11)
Verlinde’s entropic gravity reads [18]
$F=T\frac{\Delta S}{\Delta x}.$ (12)
By involving the modified equations to Verlinde’s procedure, we obtain the formula of modified emergent Newtonian gravitational force, i.e.
$\mathcal{F}=G\frac{Mm}{r^{2}}\left[1+\beta\left\{\frac{M^{2}G\hbar}{\pi r^{2}c}+\frac{\hbar^{2}}{16r^{2}}\ln\left(\frac{\pi r^{2}}{l_{p}^{2}}\right)\right\}\right].$ (13)
As we expect, if the GUP recovers to ordinary Heisenberg uncertainty principle (i.e. $\beta=0$), then Eq. (13) recovers to the ordinary familiar formula of Newtonian gravitational force, $F=G\frac{Mm}{r^{2}}$.
The equation of hydrostatic equilibrium now writes
$\frac{dP}{dr}=-\frac{GM\rho}{r^{2}}\left[1+\beta\left\{\frac{M^{2}G\hbar}{\pi r^{2}c}+\frac{\hbar^{2}}{16r^{2}}\ln\left(\frac{\pi r^{2}}{l_{p}^{2}}\right)\right\}\right],$ (14)
and for the ideal gas,
$\frac{dP}{dr}=\frac{k_{b}}{\mu\>m_{p}}\left(T\>\frac{d\rho}{dr}+\rho\>\frac{dT}{dr}\right),$ (15)
where $\mu$ is the mean molecular weight of the cluster, and $m_{p}$ is the mass of the proton. Finally, we can easily obtain the hydrostatic mass of GCs within MENG
$M_{MENG}(r)$ $=$ $-\frac{k_{b}\>T\>r}{G\>\mu\>m_{p}}\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)$ (16)
×$\times\left[1+\beta\left\{\frac{M^{2}G\hbar}{\pi r^{2}c}+\frac{\hbar^{2}}{16r^{2}}\ln\left(\frac{\pi r^{2}}{l_{p}^{2}}\right)\right\}\right].$
2.2 Hydrostatic mass of GCs within EiBI theory
Before we proceed to the derivation of GCs’ mass within EiBI theory, it is important to clarify that we involve the GUP correction on absolute temperature, that we have obtained in the subsection 2.1, in the process of the derivation. Firstly, we look at non-relativistic limit of EiBI theory which is shown by the following Poisson equation [44, 45]
$\nabla^{2}\Phi=4\pi G\rho+\frac{\kappa}{4}G\>\nabla^{2}\rho.$ (17)
Figure 1: The linear fit of the relation between $M_{MENG}$ with $\beta_{0}=2.70\times 10^{115}$ and the observed baryonic masses of 10 GCs.
Here $\Phi$ denotes gravitational potential, while $\kappa$ denotes free parameter of EiBI theory in the unit of m2. The equation of hydrostatic equilibrium within EiBI theory is given by
$\frac{dP}{dr}=-\frac{GM\rho}{r^{2}}-\frac{\kappa}{4}\>G\rho\>\frac{d\rho}{dr}.$ (18)
Now recall the equation of hydrostatic equilibrium of ideal gas shown by Eq, (15).
By combining Eq. (18) and Eq. (15), it is easy to obtain the mass of GCs within EiBI theory
$M_{EiBI}(r)=-\frac{rk_{b}}{\mu m_{p}G\rho}T\rho\left(\frac{r}{\rho}\frac{d\rho}{dr}+\frac{r}{T}\frac{dT}{dr}\right)-r^{2}\>\frac{\kappa}{4}\frac{d\rho}{dr}.$ (19)
Considering that $\left(\frac{r}{\rho}\frac{d\rho}{dr}+\frac{r}{T}\frac{dT}{dr}\right)=\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)$, we obtain
$M_{EiBI}(r)=-\frac{rk_{B}}{\mu m_{p}G}\>T\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)-\frac{r^{2}\kappa}{4}\frac{d\rho}{dr}.$ (20)
By substituting the values of the constants, we finally find the mass of GCs within EiBI theory in the unit of $\mathscr{M}_{\odot}\equiv 10^{14}\>M_{\odot}$, i.e.
$M_{EiBI}(r)=-\left[3,7\times 10^{-4}\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)+\frac{r\kappa\rho}{4M_{\odot}}\left(\frac{d\ln\rho}{d\ln r}\right)\right]\>\mathscr{M}_{\odot}.$ (21)
2.3 Hydrostatic mass of GCs within BHG
The weak-field limit of BHG is shown by the following equation [36]
$\nabla^{2}\Phi=4\pi G\rho+\frac{\Upsilon}{4}G\frac{d^{3}M}{dr^{3}},$ (22)
where $\Upsilon$ is a dimensionless BHG parameter that characterises the strength of the gravity modifications. It is worthy to note that due to higher concentration of the mass towards the center, the second term of the right-handed side of Eq. (22) corresponds to a weakening of gravity if $\Upsilon>0$ and a strengthening when the converse is true [46].
The hydrostatic equilibrium equation within BHG reads [46, 47]
$\frac{dP}{dr}=-\frac{GM\rho}{r^{2}}-\frac{\Upsilon G\rho}{4}\frac{d^{2}M}{dr^{2}},$ (23)
By combining Eq. (23) and the equation of hydrostatic equilibrium of ideal gas shown by Eq. (15), one can easily find the hydrostatic mass of GCs within BHG
$M_{BHG}=-\frac{rk_{b}T}{\mu m_{p}G}\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)-\frac{\Upsilon r^{2}}{4}\frac{d^{2}M}{dr^{2}}.$ (24)
To solve $\frac{d^{2}M}{dr^{2}}$, we use the fact that
$\frac{dM}{dr}=4\pi r^{2}\rho.$ (25)
The second derivative of $M$ with respect to $r$ gives
$\frac{d^{2}M}{dr^{2}}=4\pi r^{2}\frac{d\rho}{dr}+8\pi r\rho.$ (26)
With a little calculation, one can find
$\frac{M_{BHG}(r)}{\mathscr{M}_{\odot}}=-\frac{rk_{b}T}{\mu m_{p}G}\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)-\frac{\Upsilon\pi r^{3}\rho}{M_{\odot}}\left(2+\frac{d\ln\rho}{d\ln r}\right).$ (27)
By substituting the values of the constants, we finally find the mass of GCs within BHG in the unit of $M_{\odot}$, i.e.
$M_{BHG}=\left[{-3,7\times 10^{-4}}\left(\frac{d\ln\rho}{d\ln r}+\frac{d\ln T}{d\ln r}\right)-\frac{\Upsilon\pi r^{3}\rho}{M_{\odot}}\left(2+\frac{d\ln\rho}{d\ln r}\right)\right]\mathscr{M}_{\odot}.$ (28)
Figure 2: The linear fit of the relation between $M_{EiBI}$ with $\kappa=5.80\times 10^{40}$ and the observed baryonic masses of 10 GCs.
3 Numerical results
Vikhlinin et. al. [48] has obtained the density and temperature profiles for a total of 13 nearby relaxed galaxy clusters from Chandra and ROSAT observations. These profiles are valid up to approximately 1 Mpc [3]. We use the data of 10 GCs provided in Ref. [48] to calculate the Newtonian mass of GCs. To numerically calculate hydrostatic mass of GCs according to MENG, EiBI theory, and BHG, we use Eq. (16), Eq. (21), and Eq. (28), respectively.
Figure 3: The linear fit of the relation between $M_{BHG}$ with $\Upsilon=1.65\times 10^{71}$ and the observed baryonic masses of 10 GCs.
Table 1: Fractional difference between hydrostatic masses of GCs within various theories of gravity and Newtonian mass at $r_{500}$ for 10 clusters. The data of GCs’ mass $r_{500}$ are obtained from Ref. [3].
GC $r_{500}$ $M_{Newtonian}$ $\lvert\frac{\Delta M_{MENG}}{M_{Newtonian}}\rvert$ (%) $\lvert\frac{\Delta M_{EiBI}}{M_{Newtonian}}\rvert$ (%) $\lvert\frac{\Delta M_{BHG}}{M_{Newtonian}}\rvert$ (%)
(kpc) ($\mathscr{M}_{\odot}$) $\beta_{0}=-1.656\times 10^{110}$ $\kappa=5\times 10^{40}$ $\Upsilon=-0,1655\times 10^{69}$
A133 1007 $\pm$ 41 4.1 0.34 0.16 0.16
A383 944 $\pm$ 32 3.2 0.83 0.14 0.11
A478 1337 $\pm$ 58 7.4 0.89 0.53 0.50
A907 1096 $\pm$ 30 4.4 1.18 0.61 0.20
A1413 1299 $\pm$ 43 9.9 1.19 0.87 0.63
A1795 1235 $\pm$ 36 5.8 0.18 0.21 0.33
A1991 732 $\pm$ 33 1.4 0.56 0.53 0.03
A2029 1362 $\pm$ 43 8.6 1.08 0.75 0.60
A2390 1416 $\pm$ 48 13 0.14 0.17 0.97
MKW 4 634 $\pm$ 28 0.75 0.71 0.59 0.01
We find that for MENG, the noteworthy impact starts if we set $\beta_{0}=-1.656\times 10^{110}$; $\kappa=5\times 10^{40}$for EiBI theory, and $\Upsilon=-0,1655\times 10^{69}$ for BHG, as shown in Table 1. Even so, the impacts which are shown by the mass corrections ($\Delta M_{MENG}$, $\Delta M_{EiBI}$, and $\Delta M_{BHG}$) are so small compared to the uncertainties of Newtonian masses of GCs in Ref. [48], which are about 10%. It may be surprising if we compare our results with the results obtained in Refs. [30], [45], and [36]. In Ref. [30], $\beta_{0}$ could give a significant impact in the order of $10^{43}$; in Ref. [45], $\kappa=0,2$ could result in a meaningful impact; while in Ref. [36], $\Upsilon=0.2$ also give a noteworthy impact. It is not surprising at all if we remember that the objects used are white dwarf in Ref. [30], polytropic star in Ref. [45], and brown dwarf in Ref. [36]. These objects are really different from GCs, since their structures and EOS are really different from GCs.
Our result leads to an argument that if the values of the free parameters $\beta_{0}$, $\kappa$, and $\Upsilon$ are closer to zero than the values shown in Table 1, one can safely use the Newtonian calculation for hydrostatic mass of GCs, and the corrections of the gravitational theories used in this paper are negligible.
On the other hand, we also compare our results with the baryonic mass of GCs shown in Ref. [41]. Now consider
$\mathcal{M}=\frac{M_{modified\>gravity}}{M_{bar}},$ (29)
which represents the slope, with $M_{modified\>gravity}$ could be $M_{MENG}$, $M_{EiBI}$, or $M_{BHG}$. We expect $\mathcal{M}$ should be fundamentally unity. According to Ref. [41], significant observational uncertainties exist in the determinations of $M_{bar}$. These uncertainty also could exist in the determination of $M_{modified\>gravity}$.
For the comparison of $M_{MENG}$ with $M_{bar}$, the linear fit is shown in Fig. 1. The slope of this linear fit $0.755\pm 0.061$, which is in general agreement with the MENG in view of the presence of various uncertainties in the estimation of $M_{MENG}$ and $M_{bar}$ of the GCs. According to what we found in this work, the value of $\beta_{0}$ in this linear fit is the finest value to reach the finest slope that is close to unity.
On the other side, the linear fit of the EiBI theory with $\kappa=5.80\times 10^{40}$ gives the slope of $0.126\pm 0.086$, while the BHG with $\Upsilon=1.65\times 10^{71}$ gives the slope of $0.126\pm 0.086$. The linear fit of EiBI theory is shown by Fig. 2, while the linear fit of BHG is shown by Fig. 3. Those slope values are below the slope value of MENG, which mean that MENG is better than EiBI theory and BHG to alleviate the mass discrepancy of the hydrostatic mass and baryonic mass of GCs, with the same assumption used in Ref. [41] that there is no dark matter. It is important to note that for both EiBI theory and BHG, when we set $\kappa$ and $\Upsilon$ as positive number, the mass of GCs decrease. As we magnify both $\kappa$ and $\Upsilon$, the hydrostatic mass of GCs increases. Based on this mathematical fact, a trivial thought might arise that we should magnify the values of $\kappa$ and $\Upsilon$ to obtain the better linear fits. But in fact, it will not happen, since if we set $\kappa$ and $\Upsilon$ exceed the values used in the linear fits shown by Fig. 2 and Fig. 3, some masses of GCs reach negative values, which is not physical. By standing on this fact, we get an insight that EiBI theory and BHG cannot alleviate the mass discrepancy of hydrostatic mass and baryonic mass of GCs better than MENG.
4 Concluding remarks
In this work, we have revisited the derivation of MENG with GUP correction derived in Ref. [17]. We also derive the hydrostatic mass of GCs within EiBI theory and BHG. We applied all formulations that we have revisited and derived on 10 GCs and compared the results with the Newtonian mass obtained from Ref. [3]. The impact of the theories of gravity used in this work starts if we set $\beta_{0}=-1.656\times 10^{110}$ for MENG, $\kappa=5\times 10^{40}$ for EiBI theory, and $\Upsilon=-0.1655\times 10^{69}$ for BHG. If the values of the free parameters set closer to zero than the values mentioned before, the impact is negligible, which means that one can safely use the Newtonian formulation to calculate hydrostatic mass of GCs.
We compare our results with $M_{bar}$. The best linear fit is achieved by MENG with $\beta_{0}=2.70\times 10^{115}$, which gives the slope $\mathcal{M}$ of $0.755\pm 0.061$. This is the finest value that we found to reach the finest slope that is closest to unity. The linear fits of EiBI theory and BHG theory show much lower values of the slope compared to the MENG’s slope value. This leads us to the conclusion that in this study, MENG is the best theory of gravity that alleviates the mass discrepancy of hydrostatic mass and baryonic mass of GCs.
Acknowledgement
We are sincerely grateful to Prof. Anto Sulaksono and Idrus Husin Belfaqih for precious discussions and comments. MLP acknowledges Indonesia Endowment Fund for Education Agency (LPDP) for financial support.